You are performing a two-tailed test.

If α=.05α=.05, find the positive critical value, to three decimal places.

zα/2=zα/2=
Answer #1

Solution:

Given that,

This is two tailed test

= 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z 0.025 = 1.960

The positive critical value is ​​​​​​​1.960